Tyler’s Consulting Company has purchased a new $15,000 copier. This overhead cost will be shared by the purchasing, accounting, and information technology departments since those are the only departments which will be able to access the machine. The company has decided to allocate the cost based on the number of copies made by each department. The copier is estimated to provide 1 million copies over its life. Each division has estimated the number of copies which will be made in their department over the life of the copier.

Purchasing 350,000

Accounting 200,000

Information Tech 400,000

Note: Cost allocations are computed to 4 significant digits. How much overhead will be allocated each time a copy is made?

A. $63.3333

B. $.0158

C. $66.6667

D. $.0150
Answer #1
Answer
Correct answer is B : 0.0158
Explanation
Total number of copiers = 400000+350000+200000 =950000
Copier Rate = 15000/950000 = 0.0158
Correct answer is B : 0.0158